# Latent Gaussian dynamic factor modeling and forecasting for multivariate count time series

### v2

@@ 1 Introduction
This work develops theory, estimation, and forecasting methods for dynamic factor modeling of discrete-valued multivariate, possibly high-dimensional time series. Count time series are widespread in the natural, health, social and other sciences, for example, monthly counts of earthquakes or the amount of rainfall above a certain magnitude, daily counts of virus infections over spatial locations, item responses on surveys, and number of followers in social network over time. Mathematically, we consider a $d-$vector time series $X_{t}=(X_{i,t})_{i=1,\ldots,d}$, where $X_{i,t}\in\mathbb{N}_{0}=\{0,1,2,\ldots\}$ and $t\in\mathbb{Z}$ represents time. While the model can handle the range $\mathbb{Z}$, discrete-valued models are commonly applied to counts and the range $\mathbb{N}_{0}$. Any set of finite discrete values can be represented as a subset of $\mathbb{N}_{0}$. The primary focus will be on stationary models, though we also discuss the inclusion of covariates and differencing.
@@
In this work, we propose a multivariate, possibly high-dimensional extension of Jia et al., (2023) where the latent Gaussian series follows a DFM. A recent study by Düker et al., (2023) considered a similar extension but with the latent Gaussian series following a high-dimensional, sparse VAR model. Jia et al., (2023) considered several estimation methods, including the efficient maximum likelihood estimation based on particle approximations of likelihood function via sequential Monte Carlo. The likelihood approximation becomes computationally infeasible with available tools in higher dimensions. Both here and Düker et al., (2023), a computationally efficient scheme is employed based on a relationship between second-order properties of the observed and latent Gaussian series. The approach is also shown to have theoretical guarantees, in the high-dimensional regime $d\to\infty$. Furthermore, we suggest a novel cross-validation scheme related to model selection, namely, the number of factor series. Though we do not employ likelihood estimation based on particle approximations, we use the latter for forecasting once the model is fitted to data. This still carries an expensive computational cost but it is manageable compared to likelihood inference. Kalman filtering techniques alleviate some of the computational cost. Our modeling and forecasting approaches are illustrated in simulations and on real psychometrics data.
@@ 2.1 Model formulation
For a $d-$vector time series $\{X_{t}\}_{t\in\mathbb{Z}}=\{(X_{i,t})_{i=1,\ldots,d}\}_{t\in\mathbb{Z}}$, a latent Gaussian dynamic factor model is defined as follows. For $i=1,\ldots,d$, each component series $X_{i,t}$ is given by
$X_{i,t}=F_{i}^{-1}(\Phi(Z_{i,t}))=:G_{i}(Z_{i,t}),$ (2.1)
where $F_{i}$ is a cumulative distribution function (CDF), $F_{i}^{-1}(u)=\inf\{v:F_{i}(v)\geq{u}\}$ is its (generalized) inverse, $\Phi(z)$ is the CDF of $\mathcal{N}(0,1)$ distribution, and $Z_{i,t}$ is a zero mean, unit variance, Gaussian stationary series defined below. The CDFs $F_{i}$ are thought to come from parametric families, parameterized by a (possibly vector) parameter $\theta_{i}$. Note that by construction (2.1), $X_{i,t}$ is a stationary series and its marginal distribution is $F_{i}$. We focus here on discrete distributions $F_{i}$ taking nonnegative integer values $\mathbb{N}_{0}$. For example, if $F_{i}$ is the CDF of a Bernoulli distribution with parameter $p_{i}$, Bern$(p_{i})$, the model (2.1) becomes $X_{i,t}=\bm{1}_{\{\Phi(Z_{i,t})>1-p_{i}\}}=\bm{1}_{\{Z_{i,t}>\Phi^{-1}(1-p_{i})\}}$, where $\bm{1}_{\{\cdot\}}$ is the indicator function. More generally, if $F_{i}$ is a CDF whose support lies in $\mathbb{N}_{0}$, for example a Poisson distribution with parameter $\theta_{i}$, then $X_{i,t}$ is represented through $Z_{i,t}$ by
@@
In view of 2.2, the model discretizes the continuous-valued series $Z_{i,t}$ and is particularly natural to use in the context where $X_{i,t}$ can be thought as resulting from such discretization. In fact, (2.2) defines a count random variable $X_{i,t}$ represented through $Z_{i,t}$ that follows any marginal distribution $F_{i}$. However, we shall use the examples of Bernoulli, Poisson, negative binomial, and multinomial marginal distributions for illustration throughout the paper. For example, for multinomial marginal counts, the infinite sum in (2.2) reduces to a finite sum by the number of categories, say $N$, where $C_{i,N}=1$. Whereas the Poisson distribution with a large parameter is close to Gaussian and the latent $Z_{i,t}$’s are effectively observed, note that the Bernoulli case lies at the other extreme and is expected to be most difficult to deal with in our tasks.
We are interested in the scenario where the underlying Gaussian series $Z_{i,t}$ obeys a DFM. More specifically, we suppose that the $d-$vector time series $Z_{t}=(Z_{i,t})_{i=1,\ldots,d}$ satisfies
@@
Note that the unit variance of $Z_{i,t}$ is assumed in (2.1). For general $Z_{i,t}$, one can standardize it to have unit variance. More generally, the ACVF $\Sigma_{Z}(h)={\mathbb{E}}[Z_{t+h}Z_{t}^{\prime}]$ of $Z_{t}$ at lag $h$ can similarly become ACF $R_{Z}(h)$ as
@@ 2.2 Relation between count and Gaussian correlations
where $L_{ij}:[-1,1]\mapsto[-1,1]$ are functions to be referred to as link functions (and $L$ as a link function). Furthermore, $L_{ij}$ depends only on the CDFs $F_{i}$ and $F_{j}$ and can be expressed as described next.
@@
$L_{ij}(u)=\sum_{k=1}^{\infty}\frac{k!g_{i,k}g_{j,k}}{\Sigma_{X,ii}(0)^{1/2}\Sigma_{X,jj}(0)^{1/2}}u^{k}=:\sum_{k=1}^{\infty}\ell_{ij,k}u^{k}.$ (2.9)
@@
We will exploit (2.12) in estimation: the data for $X$ will suggest natural estimators for $L^{-1}$ and $R_{X}$, and also $R_{Z}$ by (2.12); model parameters will then be estimated from $R_{Z}$.
2.3 Nonstationary models with covariates and differencing
$\theta_{i}(t)=f(\beta_{i}^{\prime}M_{t}),$ (2.13)
where $\beta_{i}^{\prime}M_{t}$ is a linear combination of a coefficient vector and the covariate, and $f$ is a suitable function in the spirit of GLM. Then, the model (2.1) becomes
$X_{i,t}=F_{i,\theta_{i}(t)}^{-1}(\Phi(Z_{i,t}))=:G_{i,\theta_{i}(t)}(Z_{i,t}).$ (2.14)
$R_{X,ij}(t_{1},t_{2})=L_{\theta_{i}(t_{1}),\theta_{j}(t_{2})}(R_{Z,ij}(t_{1}-t_{2})),$ (2.15)
where, from (2.9),
$L_{\theta_{i}(t_{1}),\theta_{j}(t_{2})}(u)=\sum_{k=1}^{\infty}\frac{k!g_{\theta_{i}(t_{1}),k}g_{\theta_{j}(t_{2}),k}}{\textrm{Var}(X_{i,t})^{1/2}\textrm{Var}(X_{j,t})^{1/2}}u^{k}.$ (2.16)
As in (2.12), one then has
$R_{Z,ij}(h)=L_{\theta_{i}(t+h),\theta_{j}(t)}^{-1}(R_{X,ij}(t+h,t)).$ (2.17)
The estimation strategy outlined following (2.12) presents additional challenges for (2.17). While parameters $\beta_{i}$ in (2.13) can, in principle, be estimated by considering $i$th marginal data leading to $\theta_{i}(t)$, the link function $L_{\theta_{i}(t_{1}),\theta_{j}(t_{2})}$ now needs to be computed potentially for many values of $\theta_{i},\theta_{j}$. A bigger challenge is estimation of $R_{X,ij}(t+h,t)$ which now depends on $t$ and $h$. Some averaging locally over time might lead to valid estimation in a suitable framework but this is beyond the scope of this work and will be considered elsewhere.
3 Estimation
3.1 Estimation with known link function
We describe here how the parameters $\Lambda,\Psi_{1},\ldots,\Psi_{p},\Sigma_{\varepsilon}$ and $\Sigma_{\eta}$ can be estimated assuming that the marginal CDFs $F_{i}$, the link functions $L_{ij}$, $r$, and $p$ are known. (The next sections will discuss estimation and computation of $L_{ij}$, and selection of $r,p$.) Note that the relation (2.12) allows one to estimate $R_{Z}(h)=\Sigma_{Z}(h)$, $h=0,\ldots,p$, as
@@ 3.1 Estimation with known link function
Since $Z_{t}$ follows the dynamic factor model (2.3) and (2.4), the loadings matrix $\Lambda$, $\Sigma_{Y}(0)$, and $\Sigma_{\varepsilon}$ can be estimated through principal component analysis (PCA). More specifically, since the dynamic factor model (2.3) implies
@@
it is natural to estimate $\Lambda\Sigma_{Y}(0)\Lambda^{\prime}$ as an $r-$rank approximation of $R_{Z}(0)$, and take $\Sigma_{\varepsilon}$ as the approximation error. We thus proceed as follows. Consider the covariance $\widehat{R}_{Z}(0)$ estimated by (3.1). Let $\widehat{R}_{Z}(0)=\widehat{U}\widehat{E}\widehat{U}^{\prime}$ be the eigendecomposition with $\widehat{E}=\operatorname{diag}(\widehat{e}_{1},\ldots,\widehat{e}_{d})$ consisting of ordered eigenvalues $\widehat{e}_{1}\geq\ldots\geq\widehat{e}_{d}$, and $\widehat{U}=(\widehat{u}_{1},\ldots,\widehat{u}_{d})$ being the orthonormal eigenvector matrix. Setting $\widehat{U}_{r}=(\widehat{u}_{1},\ldots,\widehat{u}_{r})$ and $\widehat{E}_{r}=\operatorname{diag}(\widehat{e}_{1},\ldots,\widehat{e}_{r})$, a rank-$r$ approximation of $\Sigma_{Z}(0)$ can be taken as
@@
The relations (3.3) and (3.2) suggest setting
@@
To reiterate the idea of our approach, any estimation of the latent process (principal component analysis, Yule-Walker equations, etc.) that can be carried out on the process in terms of its second-order properties, will have its counterpart for the considered model in terms of the observable process by using relation (2.12). We shall exploit this idea again in cross-validation below (Section 3.3) when selecting $r$ and $p$. For principal component analysis, in particular, the relation (2.12) shows how the data covariance matrix $\Sigma_{X}(0)$ should be transformed entry-wise before applying the analysis.
3.2 Estimating link function and calculating its inverse
We assumed in Section 3.1 that link functions $L_{ij}$ are known. In practice, they can be estimated as follows. Recall that $L_{ij}$ is defined through marginal CDFs $F_{i}=F_{i}(\theta_{i})$ and $F_{j}=F_{j}(\theta_{j})$. The parameters $\theta_{i}$ and $\theta_{j}$ can then be estimated from the marginal distributions of the stationary data $X_{i,t}$ and $X_{j,t}$, respectively. For example, an MLE can be used for most parametric CDF models of interest. The estimator $\widehat{L}_{ij}$ can be defined through the CDFs $F_{i}(\widehat{\theta}_{i})$ and $F_{j}(\widehat{\theta}_{j})$. For example, when $F_{i}=\textrm{Bern}(p_{i})$, $\widehat{p}_{i}$ is just the sample proportion of $X_{i,t}=1$. As another example, if the $i$th marginal count series follows Poisson distribution with parameter $\theta_{i}$, it is estimated by the sample mean of observations, that is, $\widehat{\theta}_{i}=\sum_{t=1}^{T}X_{i,t}/T$. Then with the computed Hermite coefficients $\widehat{g}_{i,k}$ in (2.8), the link function for each $i$th and $j$th pair in (2.9) can be estimated by $\widehat{L}_{ij}(u)=\sum_{k=1}^{K}\widehat{\ell}_{ij,k}u^{k}$ for large enough $K$, say $K=100$, on $u\in[-1,1]$. To simplify the notation, we write $L_{ij}$ for $\widehat{L}_{ij}$ below, and are interested in calculating $L_{ij}^{-1}$.

### v3

# suppSupplementary References

@@ 1 Introduction
This work develops theory, estimation, and forecasting methods for dynamic factor modeling of discrete-valued multivariate, possibly high-dimensional time series. Count time series are widespread in the natural, health, social and other sciences, for example, monthly counts of earthquakes or the amount of rainfall above a certain magnitude, daily counts of virus infections over spatial locations, item responses on surveys, and number of followers in social network over time. Mathematically, we consider a $d-$vector time series $\{X_{t}\}_{t\in\mathbb{Z}}=\{(X_{i,t})_{i=1,\ldots,d}\}_{t\in\mathbb{Z}}$, where $X_{i,t}\in\mathbb{N}_{0}=\{0,1,2,\ldots\}$ and $t\in\mathbb{Z}$ represents time. While the model can handle the range $\mathbb{Z}$, discrete-valued models are commonly applied to counts and the range $\mathbb{N}_{0}$. Any set of finite discrete values can be represented as a subset of $\mathbb{N}_{0}$. The primary focus will be on stationary models, though we also discuss the inclusion of covariates and differencing.
@@
In this work, we propose a multivariate, possibly high-dimensional extension of Jia et al., (2023) where the latent Gaussian series follows a DFM. A recent study by Düker et al., (2024) considered a similar extension but with the latent Gaussian series following a high-dimensional, sparse VAR model. Jia et al., (2023) considered several estimation methods, including the efficient maximum likelihood estimation based on particle approximations of likelihood function via sequential Monte Carlo. The likelihood approximation becomes computationally infeasible with available tools in higher dimensions. Both here and Düker et al., (2024), a computationally efficient scheme is employed based on a relationship between second-order properties of the observed and latent Gaussian series. While Düker et al., (2024) consider a sparse VAR for the latent Gaussian series, we assume the latter to follow a DFM. DFMs are models of choice when the cross-sectional dependence across the variables is strong, for example, manifested through correlations as in our application (see Section 7 and Figure 5). Sparse VARs typically do not have this property and, in fact, there is substantial literature combining the two types of models (e.g., Lin and Michailidis, , 2020, Chen et al., , 2023). Our approach is also shown to have theoretical guarantees, in the high-dimensional regime. Here, we rely on the concentration bounds for the covariances proved in Düker et al., (2024) but substantial technical work is still needed to exploit them for the considered DFM and also to show that our factor-based model satisfies certain assumptions in Düker et al., (2024). Furthermore, we suggest a novel cross-validation scheme related to model selection, namely, the number of factor series. Though we do not employ likelihood estimation based on particle approximations, we use the latter for forecasting once the model is fitted to data. This still carries an expensive computational cost but it is manageable compared to likelihood inference. Kalman filtering techniques alleviate some of the computational cost. Our modeling and forecasting approaches are illustrated in simulations and on real psychometrics data.
@@ 2.1 Model formulation
For a $d-$vector time series $\{X_{t}\}_{t\in\mathbb{Z}}=\{(X_{i,t})_{i=1,\ldots,d}\}_{t\in\mathbb{Z}}$, a latent Gaussian dynamic factor model is defined as follows. For $i=1,\ldots,d$, each component series $X_{i,t}$ at time $t$ is given by
$X_{i,t}=F_{i}^{-1}(\Phi(Z_{i,t})):=G_{i}(Z_{i,t}),$ (2.1)
where $F_{i}$ is a cumulative distribution function (CDF), $F_{i}^{-1}(u)=\inf\{v:F_{i}(v)\geq{u}\}$ is its (generalized) inverse, $\Phi(z)$ is the CDF of $\mathcal{N}(0,1)$ distribution, and $Z_{i,t}$ is a zero mean, unit variance, Gaussian stationary series defined below. The CDFs $F_{i}$ are thought to come from parametric families, parameterized by a (possibly vector) parameter $\theta_{i}$. Note that by construction (2.1), $\{X_{i,t}\}$ is stationary and its marginal distribution is $F_{i}$. We focus here on discrete distributions $F_{i}$ taking nonnegative integer values $\mathbb{N}_{0}$. For example, if $F_{i}$ is the CDF of a Bernoulli distribution with parameter $p_{i}$, Bern$(p_{i})$, the model (2.1) becomes $X_{i,t}=\bm{1}_{\{\Phi(Z_{i,t})>1-p_{i}\}}=\bm{1}_{\{Z_{i,t}>\Phi^{-1}(1-p_{i})\}}$, where $\bm{1}_{\{\cdot\}}$ is the indicator function. More generally, if $F_{i}$ is a CDF whose support lies in $\mathbb{N}_{0}$, for example a Poisson distribution with parameter $\theta_{i}$, then $X_{i,t}$ is represented through $Z_{i,t}$ by
@@
In view of (2.2), the model discretizes the continuous-valued series $\{Z_{i,t}\}$ and is particularly natural to use in the context where $\{X_{i,t}\}$ can be thought as resulting from such discretization. In fact, (2.2) defines a count random variable $X_{i,t}$ represented through $Z_{i,t}$ that follows any marginal distribution $F_{i}$. However, we shall use the examples of Bernoulli, Poisson, negative binomial, and categorical marginal distributions for illustration throughout the paper. For example, for categorical marginal counts, the infinite sum in (2.2) reduces to a finite sum by the number of categories, say $N$, where $C_{i,N}=1$. Whereas the Poisson distribution with a large parameter is close to Gaussian and the latent $\{Z_{i,t}\}$’s are effectively observed, note that the Bernoulli case lies at the other extreme and is expected to be most difficult to deal with in our tasks.
We are interested in the scenario where the underlying Gaussian series $\{Z_{i,t}\}$ obeys a DFM. More specifically, we suppose that the $d-$vector time series $\{Z_{t}\}_{t\in\mathbb{Z}}=\{(Z_{i,t})_{i=1,\ldots,d}\}_{t\in\mathbb{Z}}$ satisfies
@@
Note that the unit variance of $Z_{i,t}$ is assumed in (2.1). For general $Z_{i,t}$, one can standardize it to have unit variance. More generally, the ACVF $\Sigma_{Z}(h)={\mathbb{E}}[Z_{t+h}Z_{t}^{\prime}]$ of $\{Z_{t}\}$ at lag $h$ can similarly become ACF $R_{Z}(h)$ as
@@ 2.2 Relation between count and Gaussian correlations
where $L_{ij}:[-1,1]\mapsto[-1,1]$ are functions to be referred to as link functions (and $L$ as a link function). Furthermore, $L_{ij}$ depends only on the CDFs $F_{i}$ and $F_{j}$ and can be expressed as described below. As $F_{i}$ depends on parameter $\theta_{i}$, we shall sometimes write $L_{\theta_{i},\theta_{j}}$ instead of $L_{ij}$ to indicate this dependence.
@@
$L_{ij}(u)=\sum_{k=1}^{\infty}\frac{k!g_{i,k}g_{j,k}}{\Sigma_{X,ii}(0)^{1/2}\Sigma_{X,jj}(0)^{1/2}}u^{k}:=\sum_{k=1}^{\infty}\ell_{ij,k}u^{k}.$ (2.9)
@@
We will exploit (2.12) in estimation in the next section.
3 Estimation of model parameters
The relation (2.12) suggests a natural estimation procedure for the model parameters of the latent series $\{Z_{t}\}$. Indeed, recall that the component function $L_{ij}$ of $L$ is defined through the marginal CDFs $F_{i}$ and $F_{j}$ which depend on the marginal parameters $\theta_{i}$ and $\theta_{j}$, respectively. As expanded upon in Section S2 of Supplemental Material, the marginal parameters $\theta_{i}$ can be estimated marginally from the stationary data $X_{i,1},\ldots,X_{i,T}$. This leads to estimates $\widehat{\theta}_{i}$ and $\widehat{\theta}_{j}$, and hence $\widehat{L}_{ij}$ and $\widehat{L}$. Computation of the inverse link functions $L_{ij}^{-1}$ (or $\widehat{L}_{ij}^{-1}$) is discussed in Section 3.2 below. On the other hand, $R_{X}(h)$ in (2.12) can be estimated through sample ACF $\widehat{R}_{X}(h)$ of the data. Substituting the sample quantities $\widehat{L}$ and $\widehat{R}_{X}(h)$ into the right-hand side of (2.12) leads to estimate $\widehat{R}_{Z}(h)$, which characterizes the second-order properties of the latent process $\{Z_{t}\}$. For the latent VAR-DFM considered in this work, the model parameters can be estimated from the model second-order properties and hence $\widehat{R}_{Z}(h)$ as described in Section 3.1. Taken together, the approach allows estimating the marginal and latent series parameters of our model.
3.1 Estimation of parameters of latent series
We describe here estimation of the parameters $\Lambda,\Psi_{1},\ldots,\Psi_{p},\Sigma_{\varepsilon}$ and $\Sigma_{\eta}$ of the latent series. We assume here that the marginal CDFs $F_{i}$ and hence the link functions $L_{ij}$ are known. In practice, the link functions can be estimated from marginal stationary data as noted above. We also assume here that $r$ and $p$ are known. Their estimation is descibed in Section 3.3 below and Section S3 of Supplemental Material. Note that the relation (2.12) allows one to estimate $R_{Z}(h)=\Sigma_{Z}(h)$, $h=0,\ldots,p$, as
@@ 3.1 Estimation of parameters of latent series
Since $\{Z_{t}\}$ follows the dynamic factor model (2.3) and (2.4), the loadings matrix $\Lambda$, $\Sigma_{Y}(0)$, and $\Sigma_{\varepsilon}$ can be estimated through principal component analysis (PCA). More specifically, since the dynamic factor model (2.3) implies
@@
it is natural to estimate $\Lambda\Sigma_{Y}(0)\Lambda^{\prime}$ as an $r-$rank approximation of $R_{Z}(0)$, and take $\Sigma_{\varepsilon}$ as the approximation error. We thus proceed as follows. Consider the covariance $\widehat{R}_{Z}(0)$ estimated by (3.1). Let $\widehat{R}_{Z}(0)=\widehat{U}\widehat{E}\widehat{U}^{\prime}$ be the eigendecomposition with $\widehat{E}=\operatorname{diag}(\widehat{e}_{1},\ldots,\widehat{e}_{d})$ consisting of ordered eigenvalues $\widehat{e}_{1}\geq\ldots\geq\widehat{e}_{d}$, and $\widehat{U}=(\widehat{u}_{1},\ldots,\widehat{u}_{d})$ being the orthonormal eigenvector matrix. Setting $\widehat{U}_{r}=(\widehat{u}_{1},\ldots,\widehat{u}_{r})$ and $\widehat{E}_{r}=\operatorname{diag}(\widehat{e}_{1},\ldots,\widehat{e}_{r})$ as $d\times r$ and $r\times r$ matrices, respectively, a rank-$r$ approximation of $\Sigma_{Z}(0)$ can be taken as
@@
The relations (3.3) and (3.2) suggest estimating the $d\times r$ loadings matrix $\Lambda$ and the $r\times r$ covariance matrix $\Sigma_{Y}(0)$ of the factor series as
@@
To reiterate the idea of our approach, any estimation of the latent process (principal component analysis, Yule-Walker equations, etc.) that can be carried out on the process in terms of its second-order properties, will have its counterpart for the considered model in terms of the observable process by using relation (2.12). We shall exploit this idea again in cross-validation below (Section 3.3) when selecting $r$. (See also Supplemental Material for choosing $p$.) For principal component analysis, in particular, the relation (2.12) shows how the data covariance matrix $\Sigma_{X}(0)$ should be transformed entry-wise before applying the analysis.
3.2 Calculating inverse of link function
In estimation, the link functions $L_{ij}$ entering $L$ in (3.1) are replaced by their sample counterparts $\widehat{L}_{ij}$. We expand here on how $\widehat{L}_{ij}$ and their inverse $\widehat{L}_{ij}^{-1}$ are computed.
Link functions $L_{ij}$ in (2.9) are defined in terms of the Hermite coefficients $g_{i,k},g_{j,k}$ which are obtained from marginal distributions $F_{i},F_{j}$ through (2.8). If $F_{i},F_{j}$ are characterized by model parameters $\theta_{i},\theta_{j}$, then $g_{i,k}=g_{i,k}(\theta_{i}),g_{j,k}=g_{j,k}(\theta_{j})$ are functions of these parameters. The parameters $\theta_{i},\theta_{j}$ can be estimated through observations for marginal $i,j$ as discussed in Section S2, leading to estimated Hermite coefficients $\widehat{g}_{i,k}=g_{i,k}(\widehat{\theta}_{i}),\widehat{g}_{j,k}=g_{j,k}(\widehat{\theta}_{j})$ and link function coefficients ^$\widehat{\ell}_{ij,k}$. Link function estimates are then $\widehat{L}_{ij}(u)=\sum_{k=1}^{K}\widehat{\ell}_{ij,k}u^{k}$ for large $K$, say $K=100$ or more. To simplify the notation, we write $L_{ij}$ for $\widehat{L}_{ij}$, and consider the computation of $L_{ij}^{-1}$ next.